# Yesnt: Are Diffusion Relighting Models Ready for Capture Stage Compositing? A Hybrid Alternative to Bridge the Gap

### v1

@@ 4 Experiments
Implementation Details
All experiments are implemented in PyTorch. We use the Adam optimizer [37] with its default parameters.
4.1 Qualitative Evaluation
[…]
To further highlight the improvements over GOF, we provide an additional comparison in Fig. 4, focusing on challenging regions such as hair. Our method recovers clean alpha values instead of relying on coarse mesh representations, resulting in more natural boundaries. We also observe smoother surface reconstruction, which is particularly noticeable in the arm region, where GOF often produces sharp edges.
Finally, we also report qualitative results on the synthetic dataset in Fig. 5. These examples show ground-truth renderings alongside our relit outputs.
GOF Mesh [77]
@@ 4.1 Qualitative Evaluation
Figure 4: Comparison of relighting quality between GOF meshes [77] (left) and our method (right). GOF often approximates natural regions such as the arm with piecewise planar surfaces, whereas our approach produces smoother reconstructions and cleaner alpha mattes in challenging areas like hair.
Figure 5: Qualitative results on the synthetic dataset, showing ground truth (top) and our method (bottom).
4.2 Numerical Analysis of Relighting Accuracy
Table 1: Quantitative comparison of relighting accuracy on synthetic data for diffusion-based methods. GOF is included for reference.
Method Conditioning PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
GOF Mesh [77] HDR Env. Map $29.4181\pm 5.7897$ $9458\pm 0.0397 0.0$$0619\pm 0.0552$
IC-Light [81] LDR Background $19.4115\pm 4.5775$ $9078\pm 0.0810 0.0$$0873\pm 0.0839$
Neural Gaffer [34] HDR+LDR Env. Map $24.4263\pm 4.5009$ $9166\pm 0.0837 0.0$$0836\pm 0.0918$
Diffusion Renderer [43] LDR+Log Env. Map $27.4926\pm 6.1947$ $9374\pm 0.0501 0.0$$0817\pm 0.0710$
Yesnt (Ours) HDR Env. Map $30.1918\pm 5.4833$ $9580\pm 0.0284 0.0$$0523\pm 0.0462$
We complement our qualitative evaluation with a quantitative analysis on object relighting on the synthetic dataset. To this end, we render ground-truth references under a wide range of environment maps and measure reconstruction accuracy across competing methods. Table 1 summarizes the results across standard image-based metrics.
Our method consistently achieves the best scores across all three metrics. In particular, it improves PSNR by a clear margin over all competing diffusion-based approaches, while also yielding the highest SSIM and lowest LPIPS. GOF serves as a strong geometric baseline, yet it is limited by hard surface boundaries and struggles in regions with complex transparency, as already seen qualitatively. In contrast, our pipeline maintains both high fidelity and perceptual quality across diverse lighting conditions.
@@ 4.3 Ablation Study
We conduct an ablation study to assess the contribution of individual components of our model. Tab. 2 reports results when removing either the relighting branch or the estimation of G-Buffer maps (roughness/metallic). Without relighting, the network reduces to a pure NVS model and achieves substantially lower accuracy. Removing the intrinsic decomposition and keeping roughness and metallic constant also results in a notable drop across all metrics, showing that physically meaningful intermediate representations are crucial for high-fidelity relighting. The full model achieves the highest scores across all metrics.
We further evaluate the effect of temporal regularization, as introduced in Sec. 3.3. Tab. 3 compares the full model against a variant without temporal regularization. The latter suffers from degraded temporal PSNR (tPSNR) and a higher perceptual path length (PPL), indicating the presence of flickering artifacts across video frames. Note, that while PPL is typically defined in the context of latent interpolations in generative models, we compute it directly across consecutive output frames to measure frame-to-frame perceptual variation. Fig. 6 illustrates this flickering effect on predicted roughness maps: the Diffusion Renderer [43] exhibits strong temporal inconsistency across consecutive frames, whereas our flow-guided smoothing produces temporally coherent results while preserving spatial detail. These findings underline the importance of temporal regularization for stable relighting in dynamic scenes.
@@
Table 3: Temporal regularization ablation on videos (see Sec. 3.3). Removing temporal regularization (w/o Temporal Reg.) leads to a decrease in temporal PSNR (tPSNR) and larger PPL, indicating stronger flickering artifacts.
@@
Figure 6: Top: Roughness maps predicted by the Diffusion Renderer [43] across three consecutive frames, showing strong temporal inconsistency. Bottom: Our flow-guided smoothing produces temporally coherent roughness estimates while preserving spatial detail.
4.4 Limitations
In terms of efficiency, the runtime of our method is dominated by the diffusion decomposition stage (cf. Tab. 4). Our method assumes a single environment‐map illumination and does not model spatially varying light sources, which limits local light transport effects (e.g., missing cast shadows on limbs). Finally, our approach relies on diffusion predictions being approximately correct on average. When the model yields inconsistent hypotheses across frames, the temporal regularizer suppresses these variations, leading to over-smoothing and a loss of fine detail and contrast.
@@ 5 Conclusion
We asked: Are diffusion relighting models ready for capture stage compositing? While they produce impressive results, they do not scale well to longer sequences. By aggregating stochastic material estimates and enforcing temporal coherence with flow-guided smoothing, our method addresses this shortcoming. Experiments on real and synthetic captures show higher fidelity and reduced flicker, bridging the gap between prototypes and production pipelines. In short, the answer is Yesnt: diffusion models alone are not yet ready, but our hybrid approach offers a practical path toward production-ready volumetric video relighting.
@@ Acknowledgments
This work has been funded by the Ministry of Culture and Science North Rhine-Westphalia under grant number PB22-063A (InVirtuo 4.0: Experimental Research in Virtual Environments). We are deeply grateful to Janelle Pfeifer for her invaluable assistance during data capture and processing.

### v2

@@ 4.1 Synthetic Analysis of Relighting Accuracy
Figure 4: Qualitative relighting on the synthetic dataset. Ground truth (top) and our method (bottom).
Figure 5: Roughness estimates show flickering in the Diff. Renderer (top) and improved temporal stability with our optimization (bottom).
Reference
@@ D. Render [33]
Figure 7: Reference (left) and our relighting (right), showing smoother surface reconstruction and cleaner alpha mattes with our approach, particularly in challenging regions such as hair.
4.2 Real World Application
Qualitative evaluation on real-world data (see Fig.˜6) indicates that prior methods face challenges in faithfully reproducing scene appearance, either by failing to model correct illumination, as in Neural Gaffer and DiffRenderer, or even by suffering from noticeable color inaccuracies, as in R3DG and IC-Light. In contrast, we observe that our method produces more faithful illumination and, relative to the relighting using a proxy mesh in Blender, yields smoother surface reconstructions, as illustrated by the arm shown in Fig.˜7, together with more realistic alpha mattes in challenging regions such as hair. The resulting increase in visual realism is further supported by a user study, reported in the supplementary material. Videos generated using our method are available on the project page.
Frame-to-Frame Coherence
Beyond static relighting quality, volumetric capture-stage compositing requires consistency across a sequence. We observe that our method maintains coherence and preserves stable material appearance, whereas diffusion-only approaches frequently exhibit temporal inconsistency, particularly when constrained to short windows imposed by memory limitations (see also Fig.˜5). These observations are supported by a user study (see supplementary material), in which participants rated our results as significantly more temporally stable, with a very large effect size for temporal smoothness (Cohen’s $d=1.57$).
@@ 4.3 Ablation Study
We conduct an ablation study to assess the contribution of individual components of our model. Tab.˜2 reports results when removing either the relighting branch or the estimation of G-Buffer maps (roughness/metallic). Without relighting, the network reduces to a pure NVS model and achieves substantially lower accuracy. Removing the intrinsic decomposition and keeping roughness and metallic constant also results in a notable drop across all metrics, showing that physically meaningful intermediate representations are crucial for high-fidelity relighting. The full model achieves the highest scores across all metrics.
Table 3: Temporal regularization ablation on videos (see Sec.˜3.3). Removing temporal regularization (w/o Temporal Reg.) leads to a decrease in temporal PSNR (tPSNR) and larger PPL, indicating stronger flickering artifacts.
@@
We further evaluate the effect of temporal regularization, as introduced in Sec.˜3.3. Tab.˜3 compares the full model against a variant without temporal regularization. The latter suffers from degraded temporal PSNR (tPSNR) and a higher perceptual path length (PPL), indicating the presence of flickering artifacts across video frames. Note, that while PPL is typically defined in the context of latent interpolations in generative models, we compute it directly across consecutive output frames to measure frame-to-frame perceptual variation. Fig.˜5 illustrates this flickering effect on predicted roughness maps: the Diffusion Renderer [33] exhibits strong temporal inconsistency across consecutive frames, whereas our flow-guided smoothing produces temporally coherent results while preserving spatial detail. These findings underline the importance of temporal regularization for stable relighting in dynamic scenes.
4.4 Limitations
@@
The runtime of our method scales proportionally with sequence length, with the diffusion-based decomposition forming the primary computational bottleneck and limiting throughput for long sequences (cf. Tab.˜4). While this currently restricts the method to offline processing, it remains compatible with typical volumetric capture workflows as its computational overhead is negligible compared to GOF’s optimization cost.
Furthermore, our method assumes a single environment‐map illumination and does not model spatially varying light sources, which limits local light transport effects (e.g., missing cast shadows on limbs).
Finally, our method relies on the assumption that diffusion-based material estimates are accurate on average. While this holds in most cases, rare failures occur when the diffusion model produces strongly inconsistent predictions across frames. In these cases, temporal regularization may suppress such outliers, leading to over-smoothing of fine-scale detail.
@@ 5 Conclusion
We asked: Are diffusion relighting models ready for capture stage compositing? While they produce impressive results, they do not scale well to longer sequences. By aggregating stochastic material estimates and enforcing temporal coherence with flow-guided smoothing, our method addresses this shortcoming. Experiments on real and synthetic captures show higher fidelity and reduced flickering, bridging the gap between prototypes and production pipelines. Our results suggest that enforcing structured temporal constraints is more effective than increasing generative model capacity for achieving stable video relighting. In short, the answer is Yesnt: diffusion models alone are not yet ready, but our hybrid approach offers a practical path toward production-ready volumetric video relighting.
@@ Acknowledgments
This work has been funded by the Ministry of Culture and Science North Rhine-Westphalia under grant number PB22-063A (InVirtuo 4.0: Experimental Research in Virtual Environments).